OpenStudy (khally92):

Find an equation in simplified form for the set of all points whose distance to (1,4,6) is twice the distance to (2,4,0) . Name and give a precise description of this surface Help Please.

ganeshie8 (ganeshie8):

Let \(P=(x,y,z)\) be any point on the required surface

ganeshie8 (ganeshie8):

remember the distance formula in 3D ?

OpenStudy (khally92):

yes

OpenStudy (khally92):

I KNOW THE DISTANCE FORMUL IN 3D

ganeshie8 (ganeshie8):

Let M = (1,4,6) and N = (2,4,0) .

ganeshie8 (ganeshie8):

PM = 2*PN

ganeshie8 (ganeshie8):

use distance formula

OpenStudy (khally92):

you mean {(1-x )^2+( 4-y)^2 + (6-z)^2}^1/2=2PN

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

squaring both sides gets rid of radicals : \[(x-1)^2+(y-4)^2+(z-6)^2 = 4[(x-2)^2+(y-4)^2+(z-0)^2]\]

ganeshie8 (ganeshie8):

try simplifying

OpenStudy (khally92):

Ok thank you, I got 2x-12z+33=0 I will check my work again thank you so much really appreciate.

ganeshie8 (ganeshie8):

you should get a sphere

OpenStudy (khally92):

Awesome thank you, forgot to multiplied the right side by 4. thank you.

ganeshie8 (ganeshie8):

np :)